OpenStudy (anonymous):

Fan and Medal, Please Help Samuel hit the baseball from 2 feet off the ground at an initial vertical velocity of 30 feet per second. How long will it take the ball to hit the ground? A. Write the formula. B. Solve the problem. C. State your answer. Round your answer to the nearest hundredth of a second.

OpenStudy (sweetburger):

Wouldn't your physics text book have almost exactly how do this in it.

OpenStudy (anonymous):

lol this isn't physics, its Algebra @sweetburger

OpenStudy (anonymous):

34/102 = BP #1 24/96 = BP #2 34/102 = 1/3 24/96 = 1/4 1/3 --> 4/12 1/4 --> 3/12 4/12 + 3/12 = 7/12 7/12

OpenStudy (anonymous):

@tkhunny help please

OpenStudy (tkhunny):

Standard stuff, given the right equation. Usually, \(y(t) = -16t^{2} + v_{0}t + h_{0}\) \(-16\) represents the acceleration due to gravity. \(v_{0}\) represents the initial velocity. \(h_{0}\) represents the initial height. You are given: \(y(t) = -16t^{2} + 30t + 2\) Your task is to solve for t such that \(y(t) = 0\) It's just a quadratic equation. Do it!!

OpenStudy (anonymous):

Can I factor by grouping or do I have to use the quadratic formula?

OpenStudy (tkhunny):

Who cares? Whatever works. Note: "Factor by Grouping" is a phantom. It is rarely useful. Once in a while it pops out at you, so keep your mind open to it. Generally, though, sing it a lullaby.